# SLA Management in Reconfigurable Multi-Agent RAG: A Systems Approach to Question Answering

### v1

@@ V-A Preprocessing Methods
  • Re-ranking: Utilized a cross-encoder model based on ’ms-marco-MiniLM-L-12-v2’ from the Sentence Transformers library [21]. The cross-encoder accepts the query and document as input and returns a relevance score between 0 and 1. Documents are re-ranked based on relevance scores before thresholding.  [^5] Note: The raw entity JSON is not used directly as input. Search configurations define an LLM-friendly view of the entity, specifying which fields should be included. The design of this view generation model is outside the scope of this work.
V-B Arbitration
@@
V-C Algorithm Ablations
To evaluate the effectiveness of our arbitration algorithm and understand the impact of its components, we conducted a series of ablation experiments:
  • Random Candidate Selection:
    • Method: Select a candidate answer at random from the set of affirmative responses, bypassing the arbitration function.
    • Purpose: Establish a baseline to assess the added value of the arbitration function in selecting higher-quality answers compared to random selection.
  • Voting Procedure with Random Positive Selection:
    • Method: Apply the threshold evaluation to determine if sufficient affirmative responses exist. If so, randomly select one of the affirmative answers without applying $A$.
    • Purpose: Isolate the effect of the threshold-based decision while assessing the necessity of the arbitration function for answer selection.
  • Cross-Encoder Arbitration on All Candidate Responses:
    • Method: Apply the cross-encoder-based arbitration function $A$ to all candidate responses, including both affirmative and negative ones (e.g., null or placeholder answers).
    • Purpose: Examine whether including all candidate responses in the arbitration process affects the quality of the final answer and whether the arbitration function can effectively discriminate among all types of responses.
These ablation experiments allowed us to analyze the contributions of different components of the arbitration algorithm:
  • Effectiveness of the Arbitration Function $A$:
    • By comparing the random selection methods with the cross-encoder arbitration, we assessed how much the arbitration function improves answer quality by selecting the most relevant response.
  • Impact of the Decision Threshold $T$:
    • Evaluating different values of $T$ helped determine the optimal balance between answer availability and quality.
    • A higher threshold requires more consensus among agents before providing an answer, potentially reducing incorrect or low-quality responses.
  • Inclusion of Negative Responses:
    • Testing the arbitration function on all candidate responses explored whether the model could effectively handle and filter out non-informative or irrelevant answers.
VI Results
@@
VI-A Single-Agent Experiments
The results of the single-agent experiments are summarized in the upper section of Table I. These results encompass the three preprocessing strategies discussed in the Single Agent Experiments section
@@ VI-A Single-Agent Experiments
TABLE I: Experimental Results for Single and Multi Agent Experiments
@@
VI-B Multi-Agent Experiments
The multi-agent experiments were designed to evaluate the impact of ensemble configurations on system performance. We varied the number of agents in the ensemble and observed the effects of different arbitration mechanisms, all of which used a threshold $T=0.5$, requiring at least half of the agents to provide an affirmative answer for a response to be returned. The results can be found in the bottom section of Table I. Detailed charts are shown in Figures 6 - 10.
@@ VI-B Multi-Agent Experiments
Figure 6: Recall vs. Ensemble Size
Figure 7: Precision vs. Ensemble Size
Figure 8: F1 vs. Ensemble Size
Figure 9: Hallucination Rate vs. Ensemble Size
Figure 10: Incongruent Response Rate vs. Ensemble Size
VII Discussion
Our experiments demonstrate the effectiveness of dynamically reconfigurable multi-agent RAG systems in managing trade-offs between answer quality and resource utilization. Insights Effectiveness of Aggressive Thresholding: By aggressively limiting the context to the most relevant information, we reduce noise and enhance the LLM’s ability to generate accurate answers. Limitations of Re-ranking Alone: While re-ranking aims to prioritize relevant documents, it can inadvertently discard essential context, adversely affecting performance. Scalability of Multi-Agent Systems: Adding more agents generally improves performance but increases cost and latency. Systems must be designed to adjust dynamically based on SLA requirements.
VII-A Future Work
  • Exploration of Additional Arbitration Mechanisms: Investigate other methods, such as weighted voting or machine learning-based arbitration, to enhance decision-making.
  • Incorporation of QoS Parameters: Extend the framework to include latency and computational constraints explicitly, enabling more granular SLA management.
  • Complexity of Planning Mechanisms: Develop more sophisticated Planning Modules capable of temporal replication (e.g., handling requests for refinement) and adjusting strategies in real-time.
  • Simulation of Real-World Conditions: Incorporate factors like network congestion and varying computational capacities to assess the system’s robustness under diverse operational environments.
  • Threshold Parameter Optimization: Study the impact of different threshold values in arbitration mechanisms to find optimal settings for various use cases.
Our case study illustrates that a systems-oriented approach to SLA management in reconfigurable multi-agent RAG systems is both feasible and beneficial. By dynamically adjusting system parameters based on non-functional requirements, we can effectively balance answer quality with resource constraints. This work serves as a foundational step toward deploying RAG-based QA systems in real-world applications with diverse and stringent SLA demands.

### v2

@@ 6.1. Preprocessing Methods
  • Re-ranking: Utilized a cross-encoder model based on ‘ms-marco-MiniLM-L-12-v2’ from the Sentence Transformers library (Reimers and Gurevych, 2019). The cross-encoder accepts the query and document as input and returns a relevance score between 0 and 1. Documents are re-ranked based on relevance scores before thresholding. [^7] Note: The raw entity JSON is not used directly as input. Search configurations define an LLM-friendly view of the entity, specifying which fields should be included. The design of this view generation model is outside the scope of this work.
6.2. Arbitration
@@
6.3. Algorithm Ablations
To isolate the contributions of each component in our arbitration algorithm, we conducted a series of ablation experiments. First, we bypassed the arbitration function entirely by selecting a candidate answer at random from the set of affirmative responses. Next, we evaluated the effect of the threshold mechanism by randomly choosing a response from the affirmative set after threshold evaluation. Finally, we applied the cross-encoder-based arbitration function $A$ over the complete set of candidate responses to assess its overall impact on the final output.
7. Results
@@
7.1. Single-Agent Experiments
The results of the single-agent experiments are summarized in the upper section of Table 2. These results encompass the three preprocessing strategies discussed in the Single Agent Experiments section
@@ 7.1. Single-Agent Experiments
Table 2. Experimental Results for Single and Multi-Agent Experiments
@@
7.2. Multi-Agent Experiments
The multi-agent experiments were designed to evaluate the impact of ensemble configurations on system performance. We varied the number of agents in the ensemble and observed the effects of different arbitration mechanisms, all of which used a threshold $T=0.5$, requiring at least half of the agents to provide an affirmative answer for a response to be returned. The results can be found in the bottom section of Table 2. Detailed charts are shown in Figures 6 - 10.
@@ 7.2. Multi-Agent Experiments
Figure 6. Recall vs. Ensemble Size
Figure 7. Precision vs. Ensemble Size
Figure 8. F1 vs. Ensemble Size
Figure 9. Hallucination Rate vs. Ensemble Size
Figure 10. Incongruent Response Rate vs. Ensemble Size
8. Discussion and Future Directions
In this work, we introduce a dynamically reconfigurable multi-agent RAG system and present experimental results that highlight its ability to navigate multiple SLA dimensions—balancing answer quality and resource efficiency. Our results show that aggressive thresholding at the retrieval stage, by focusing on the most pertinent context, enables the system to generate more responses with improved precision. However, conventional re-ranking methods, while effective at prioritizing relevant documents, can sometimes filter out essential context and thereby impact the overall performance balance. Similarly, scaling the number of agents tends to boost metrics like precision and recall, but this comes with increased resource consumption. These interdependent trade-offs make the case for a system that can adapt in real time to diverse SLA requirements.
Looking ahead, several promising research directions are apparent. Exploring alternative arbitration mechanisms—such as weighted voting or additional machine learning-based approaches—may help optimize overall system performance, even if improvements in one dimension might be offset by challenges in another. Incorporating additional explicit QoS parameters—particularly latency and computational constraints—into our experiments would enable a more nuanced approach to SLA management. Coupling this with simulations of real-world conditions, such as network congestion and variable computational resources, and optimizing threshold parameters will be critical for comprehensively understanding their impact on latency and other performance dimensions. We would also like to explore enhancements to our planning modules that enable temporal replication and dynamic strategy adjustments, which could further improve the system’s overall adaptability.
Overall, our work demonstrates that a systems-oriented approach to SLA management in reconfigurable multi-agent RAG systems is not only feasible but also essential for deploying QA solutions that meet the multifaceted demands of real-world applications.